OpenStudy (anonymous):

4) The DIAMETER of a sphere is doubled. What is the new volume of the sphere? A) 8 3 πr3 B) 16 3 πr3 C) 32 3 πr3 D) 64 3

OpenStudy (anonymous):

C .

OpenStudy (anonymous):

i hav justifications if u want i'll tell u

OpenStudy (anonymous):

Why Do You Say Its C ? lol

OpenStudy (anonymous):

When you double the diameter, you double the radius. V = (4/3)(pi)r^3, so new V is (4/3)(pi)(2r)^3, so the volume gets to be 8 times as much as the original.

OpenStudy (anonymous):

let the dia be 'd' the ques says that the dia is doubled therefore the new dia is '2d' => radius of the new sphere is '2r' we know that, volume of spere is v,= (4/3) pie (radius)^3 = (4/3) pie (2r)^3 if you solve it u'll be getting ma ans

OpenStudy (anonymous):

(4/3)(pi)(2r)^3 = (4/3)(pi)(2^3)(r^3) = (4/3)(pi)8(r^3) = 8(4/3)(pi)(r^3) = 8 times original volume.

OpenStudy (anonymous):

Is this making sense to you now?

OpenStudy (anonymous):

Sorry , Umm Yea Im Starting To Get A Better Understanding